At a tennis court they charge 10.00 for feeding the ball 16 times. At this rate, how much will 120 times of ball feeding cost?
Answer by checkley79(3341)

10/16=X/120
16X=10*120
16X=1200
X=1200/16
X=$75 IS THE CHARGE.